OpenStudy (anonymous):

Please Help use synthetic division and the given factor to completely factor the polynomial. x^3-x^2-24x-36;(x-6) Show all work

OpenStudy (ranga):

Do you know synthetic division? YouTube has several examples.

OpenStudy (anonymous):

no can you help walk me through the problem?

OpenStudy (ranga):

Rather than explain by typing I will see if I can find a good video:

OpenStudy (ranga):

http://www.youtube.com/watch?v=bZoMz1Cy1T4

OpenStudy (anonymous):

can you just try to explain? my computer cant handle videos all too well

OpenStudy (ranga):

Divide x^3-x^2-24x-36 by (x-6) using synthetic division. Write the coefficients of the numerator: 1 -1 -24 -36 To divide by x - 6 take the opposite sign of -6. That will be +6 ________________________ 6 | 1 -1 -24 -36 | 6 +30 +36 |_____________________________ 1 5 6 0 Last line is quotient: x^2 + 5x + 6 Remainder = 0 x^3-x^2-24x-36 = (x-6)(x^2 + 5x + 6) x^2 + 5x + 6 is a quadratic expression which you can factor.